Add these two fractions, after converting them to their LCD: 13/18 + 8/27?
I got the 13/18 but I can get the 8/27 I just don't get it. Thank you for your help in advance.

Found 2 solutions by Alan3354, Jessa24:
Answer by Alan3354(69443)

Add these two fractions, after converting them to their LCD: 13/18 + 8/27?
I got the 13/18 but I can get the 8/27 I just don't get it. Thank you for your help in advance.
-------------
I don't know what "I got the 13/18 " means ??
-------------
13/18 + 8/27
A common denominator is always the product = 27*18 = 486
13/18 + 8/27 = 351/486 + 144/486 = 495/486
= 55/54
-----------
Using the LCD, 54:
13/18 + 8/27 = 39/54 + 16/54 = 55/54 same answer.
Don't waste time on a test finding the LCD, any common denominator works.

Answer by Jessa24(1)

Sorry that was not to be there.